.  y”  -  y’ -   2y  =  1 – 2x - 9e-x        


1
Expert's answer
2021-02-24T06:52:04-0500

Solution:

We know that the general solution will be the sum of the complementary solution and particular solution.

First we find the complementary solution by solving "\frac{d^{2} y(x)}{d x^{2}}-\frac{d y(x)}{d x}-2 y(x)=0" :

Assume a solution will be proportional to "e^{\lambda x}" for some constant "\lambda" .

Substitute "y(x)=e^{\lambda x}" into the differential equation:

"\frac{d^{2}}{d x^{2}}\left(e^{\lambda x}\right)-\frac{d}{d x}\left(e^{\lambda x}\right)-2 e^{\lambda x}=0"

"\lambda^{2} e^{\lambda x}-\lambda e^{\lambda x}-2 e^{\lambda x}=0"

"\left(\lambda^{2}-\lambda-2\right) e^{\lambda x}=0"

Since "e^{\lambda x} \neq 0" for any finite "\lambda" , the zeros must come from the polynomial:

"\lambda^{2}-\lambda-2=0"

"(\lambda-2)(\lambda+1)=0"

"\lambda=-1 \ or\ \lambda=2"

The general solution is the sum of the above solutions:

"y(x)=y_{1}(x)+y_{2}(x)=c_{1} e^{-x}+c_{2} e^{2 x}"

Now we determine the particular solution by the method of undetermined coefficients:

The particular solution will be the sum of the particular solutions to "\frac{d^{2} y(x)}{d x^{2}}-\frac{d y(x)}{d x}-2 y(x)=-2 x+1\ and\ \frac{d^{2} y(x)}{d x^{2}}-\frac{d y(x)}{d x}-2 y(x)=-9 e^{-x}"

The particular solution to "\frac{d^{2} y(x)}{d x^{2}}-\frac{d y(x)}{d x}-2 y(x)=-2 x+1" is of the form:

"y_{p_{1}}(x)=a_{1}+a_{2} x"

The particular solution to "\frac{d^{2} y(x)}{d x^{2}}-\frac{d y(x)}{d x}-2 y(x)=-9 e^{-x}" is of the form:

"y_{p_{2}}(x)=x\left(a_{3} e^{-x}\right)" , where "a_{3} e^{-x}" was multiplied by "x" to account for "e^{-x}" in the complementary solution.

"y_{p}(x)=y_{p_{1}}(x)+y_{p_{2}}(x)=a_{1}+a_{2} x+a_{3} e^{-x} x"

Now, we solve for the unknown constants "a_{1}, a_{2},\ and\ a_{3}" :

"\begin{array}{l}
\text { Compute } \frac{d y_{p}(x)}{d x}: \\
\frac{d y_{p}(x)}{d x}=\frac{d}{d x}\left(a_{1}+a_{2} x+a_{3} e^{-x} x\right) \\
=a_{2}+a_{3} e^{-x}-a_{3} e^{-x} x \\
\text { Compute } \frac{d^{2} y_{p}(x)}{d x^{2}}: \\
\frac{d^{2} y_{p}(x)}{d x^{2}}=\frac{d^{2}}{d x^{2}}\left(a_{1}+a_{2} x+a_{3} e^{-x} x\right) \\
=a_{3}\left(-2 e^{-x}+e^{-x} x\right)
\end{array}"

Substitute the particular solution "y_{p}(x)" into the differential equation:

"\frac{d^{2} y_{p}(x)}{d x^{2}}-\frac{d y_{p}(x)}{d x}-2 y_{p}(x)=-9 e^{-x}-2 x+1"

"\begin{array}{l}

a_{3}\left(-2 e^{-x}+e^{-x} x\right)-\left(a_{2}+a_{3} e^{-x}-a_{3} e^{-x} x\right)-2\left(a_{1}+a_{2} x+a_{3} e^{-x} x\right) \\

\quad=-9 e^{-x}-2 x+1

\end{array}"

"-2 a_{1}-a_{2}-3 a_{3} e^{-x}-2 a_{2} x=1-9 e^{-x}-2 x"

On equating the coefficients on both sides of the equation:

"-2 a_{1}-a_{2}=1"

"-3 a_{3}=-9"

"-2 a_{2}=-2"

On solving,

"\begin{array}{l}

a_{1}=-1 \\

a_{2}=1 \\

a_{3}=3

\end{array}"

Substitute

"a_{1}, a_{2},\ and\ a_{3}\ into\ y_{p}(x)=a_{1}+x a_{2}+e^{-x} x a_{3}
\\ We\ get\ 
\\
y_{p}(x)=x+3 e^{-x} x-1"

So, the general solution is: "y(x)=y_{c}(x)+y_{p}(x)=x+3 e^{-x} x+c_{1} e^{-x}+c_{2} e^{2 x}-1"